You own a lot in Montreal that is currently unused. Similar lots have recently sold for $1.9 million. Over the past five years, the price of land in the area has increased 12 percent per year, with an annual standard deviation of 25 percent. A buyer has recently approached you and wants an option to buy the land in the next 12 months for $2.1 million. The risk-free rate of interest is 5 percent per year, compounded continuously. Suppose you wanted the option to sell the land to the buyer in one year. What is the price of the transaction today?
Answer #1

According to Black - schole Option pricing model the Option price will be as follows:-

Formulas used:-

ln(S0/K)=IFERROR(LN(C5/C6),"na")
(r+σ2/2)t=(C8+(C9^2)/2)*C7
σ√t=C9*SQRT(C7)
d1=IFERROR((C13+C14)/C15,"na")
d2=IFERROR(C16-C15,"na")
N(d1)=IFERROR(NORM.S.DIST(C16,TRUE),"na")
N(d2)=IFERROR(NORM.S.DIST(C17,TRUE),"na")
N(-d1)=IFERROR(NORM.S.DIST(-C16,TRUE),"na")
N(-d2)=IFERROR(NORM.S.DIST(-C17,TRUE),"na")
e-rt =EXP(-C8*C7)

Option Price=IFERROR(IF(C4='--> Additional Info'!A3,C5*C18-C6*C22*C19,IF(C4='--> Additional Info'!A4,C6*C22*C21-C5*C20,"na")),"na")

I hope my efforts will be fruitful to you...?